OpenStudy (anonymous):

Medal&Fan Solve (9x + 2) – 3(x – 4) = 6. Select one: a. x = 4/3 b. x = –4/3 c. It is an identity and is true for all real numbers. d. There is no solution.

OpenStudy (nikato):

Can you distribute the -3(x-4)?

OpenStudy (anonymous):

yeahh. I think so.

OpenStudy (nikato):

Ok. So what is your equation now after distributing?

OpenStudy (anonymous):

9x +2 -3x - 12 =6

OpenStudy (nikato):

Ok. Correct. And now can you combine like terms on the left side?

OpenStudy (anonymous):

6x -10 = 6

OpenStudy (nikato):

Yes. And now add 10 to both sides. So you're left with only the variable term on one side

OpenStudy (anonymous):

6x=16

OpenStudy (nikato):

Correct. Now divide both sides by 6 so that you will have x=?

OpenStudy (anonymous):

soo its D?

OpenStudy (nikato):

Wait, hold up. I think something's not right

OpenStudy (anonymous):

I mean its not A or B, so C or D

OpenStudy (nikato):

My bad. You actually distributed wrong. Should be 9x+2-3x+12=6. It's +12 not -12 Try to solve for x now

OpenStudy (nikato):

We're you able to find x?